# Computing an approximation of the partial Weyl closure of a holonomic module

### v1

@@ Abstract.
The Weyl closure is a basic operation in algebraic analysis: it converts a system of differential operators with rational coefficients into an equivalent system with polynomial coefficients. In addition to encoding finer information on the singularities of the system, it serves as a preparatory step for many algorithms in symbolic integration. A new algorithm is introduced to approximate the partial Weyl closure of a holonomic module, where the closure is taken with respect to a subset of the variables. The method is based on a non-commutative generalization of Rabinowitsch’s trick and yields a holonomic module included in the Weyl closure of the input system. The algorithm is implemented in the Julia package MultivariateCreativeTelescoping.jl and shows substantial speedups over existing exact Weyl closure algorithms in Singular and Macaulay2.
Weyl closure, holonomic modules, Weyl algebras, Gröbner bases, symbolic integration
@@ 1. Introduction
A recent integration algorithm (Brochet et al., 2025) assumes that the parameters $\mathbf{t}=(t_{1},\dots,t_{m})$ of the integrals are rational variable, while the integration variables $\mathbf{x}$ are polynomial variables, and takes as input generators of $\operatorname{ann}_{W_{t,\mathbf{x}}(\mathbf{t})}(f)$ where $W_{t,\mathbf{x}}(\mathbf{t})=\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\partial_{\mathbf{t}},\partial_{\mathbf{x}}\rangle$. This motivates extending the definition of Weyl closure to partial Weyl closure. Let $S\subseteq W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}$ be a left submodule. The partial Weyl closure of $S$ with respect to $\mathbf{x}$ is
@@
The problem of computing generators of $\operatorname{Cl}(S)$ given $S\subset W_{\mathbf{x}}^{r}$ was first introduced by Chyzak and Salvy (Chyzak and Salvy, 1998) for the case $r=1$, under the name of extension/contraction problem. This problem was later solved by Tsai for general $r$ in his PhD thesis (Tsai, 2000b, a). He showed (Tsai, 2000a, Theorem 2.2.1) that the Weyl closure of $S\subset W_{\mathbf{x}}^{r}$ can be computed by inverting a single polynomial $p$ vanishing on the singular locus of $S$:
@@
The computation of annihilators of holonomic functions in the rational Weyl algebra $W_{\mathbf{x}}(\mathbf{x})$ is already well understood. Systems of differential equations are known for basic functions and algorithmic closure by sum, product and more (Comtet, 1964; Takayama, 1992) exist for deducing systems for more complicated ones. These algorithms are implemented in Maple (Chyzak, ), Mathematica (Koutschan, 2014), and Sage (Kauers et al., 2015). Although similar closure properties for sum and product exist in the Weyl algebra (Oaku and Takayama, 2001) (see also (Oaku, 2013, Sec.3)), their computational efficiency seem unclear. For example, the closure by product requires to double the number of variables and eliminate half of them by Gröbner basis computation, which is known to be costly.
@@
The notion of holonomicity is extended to $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-modules and a new algorithm is proposed for computing an holonomic approximation of the partial Weyl closure. That is, given a finite rank module $S\subset W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}$, the algorithm computes a module $S^{\prime}$ such that $S^{\prime}\subseteq\operatorname{Cl}(S)$ and $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}/S^{\prime}$ is holonomic. The algorithm proceeds by iteratively constructing a sequence of modules
@@
The algorithm is based on a generalization of Rabinowitsch’s trick to D-modules, which consists in introducing a new variables $T$ thought of as $1/p$ for some $p\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$, together with the additional equation pT1=0pT-1=0. The commutation rule between $T$ and $\partial_{\mathbf{x}},\partial_{\mathbf{t}}$ are not well suited for computing Gröbner basis in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ for an order eliminating $T$. This problem is tackled by computing Gröbner bases in $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})[T]^{r}$ seen as an infinite-rank $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-module (Takayama, 1990). Then, saturations of $S$ with respect to $p$ are computed iteratively until the stopping condition is reached. Lastly, the algorithm has been implemented in the Julia package MultivariateCreativeTelescoping.jl (Brochet, 2025b). It compares very favorably to existing implementation, both because of a better Gröbner basis implementation and because of an algorithmic improvement.
Compared with Tsai’s approach, the present algorithm avoids several computationally expensive steps. Tsai’s algorithm first computes generators of the $W_{\mathbf{x}}$-module $W_{\mathbf{x}}[1/p]S$ using the localization algorithm of Oaku, Takayama, and Walther (Oaku et al., 2000), and then recovers the Weyl closure by performing the intersection $W_{\mathbf{x}}[1/p]S\cap W_{\mathbf{x}}^{r}$. The localization step requires the computation of a $b$-function associated with a weight vector of a suitable ideal, which is known to be difficult in general. Moreover, simplifications with respect to the polynomial $p$ are postponed until the final intersection step. In particular, when $S$ is an ideal annihilating a function $f$, the localization algorithm returns an ideal $S^{\prime}$ together with an integer $\ell$ such that $S^{\prime}$ annihilates $f/p^{\ell}$, which typically leads to a significant blow-up in the size of expressions compared to $S$ and $\operatorname{Cl}(S)$. This issue is avoided in the present algorithm by performing simplifications with respect to $p$ as early as possible through iterative saturation.
@@ 2.2. Singular Locus
The principal symbol (or initial form) of $P$ with respect to $(\bm{0},\bm{1})$ is the element in $\mathbb{K}[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$ defined by
@@
$\operatorname{in}_{(\bm{0},\bm{1})}(S)\subset\mathbb{K}[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$
@@
$\operatorname{char}(M)=V\!\left(\operatorname{in}_{(\bm{0},\bm{1})}(S)\right)\;\subset\;\mathbb{C}^{2n+m}.$
@@
$\mathbb{C}^{2n+m}\to\mathbb{C}^{n},\qquad(\mathbf{x},{\bm{\xi}},{\bm{\zeta}})\mapsto\mathbf{x}.$
@@
$\operatorname{Sing}(I)=V\left(\left(\operatorname{in}_{(\bm{0},\bm{1})}(S):\langle\xi_{1},\dots,\xi_{n},\zeta_{1},\dots,\zeta_{m}\rangle^{\infty}\right)\cap\mathbb{C}[\mathbf{x}]\right),$
where, for a module $A\subset\mathbb{K}[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$, and an ideal $B\subset\mathbb{K}[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]$ the saturation of $A$ with respect to $B$ is given by
(10) $A:B^{\infty}=\left\{a\in\mathbb{K}[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}\;\middle|\;\exists\,k\in\mathbb{N}\text{ such that }B^{k}a\subset A\right\}.$
@@ 3. Holonomic $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-modules
When $m=0$ this definition coincides with the standard definition of holonomic $W_{\mathbf{x}}$-modules by Theorem 1, and when $n=0$ it coincides with the definition of D-finite modules, that is, modules that are finite-dimensional over $\mathbb{K}(\mathbf{t})$. The equivalence between the second and third point of Theorem 1, which gave an algorithmic criterion for testing holonomicity, naturally extends to this setting.
@@ 4. Partial Weyl closure by saturation
The goal of this section is to prove Theorem 1, which states that the partial Weyl closure of a module $S\subset W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ can be computed by inverting a single polynomial vanishing on the singular locus of $S$. A similar result was proved by Tsai (Tsai, 2000a, Theorem 2.2.1) in the case where no parameter $\mathbf{t}$ is present. The argument given here is a straightforward extension of his approach to this mixed setting.
@@
Let $S$ be a submodule of $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ and let $f\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ be a polynomial vanishing on $\operatorname{Sing}(S)$. Then,
@@
Proof.
Proof of Lemma 2. Let $T$ be the torsion module
@@
annihilates every element of $T$. Applying any $\delta\in\{\bm{\partial}_{\mathbf{x}},\bm{\partial}_{\mathbf{t}}\}$ gives
@@
Indeed, since $S$ is finite-rank module, for any $i\in\llbracket 1,r\rrbracket$ and any $\delta\in\{\bm{\partial}_{\mathbf{x}},\bm{\partial}_{\mathbf{t}}\}$, there exists a nonzero element $L^{\prime}\in W_{\mathbf{t},\mathbf{x}}(\mathbf{t},\mathbf{x})S\cap\mathbb{K}(\mathbf{t},\mathbf{x})\langle\delta\rangle e_{i}$. Multiplying it by a suitable polynomial, we obtain a non-zero element $L\in S\cap\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\delta\rangle e_{i}$. Its initial form $\operatorname{in}_{(\bm{0},\bm{1})}(L)$ then belongs to $\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\delta\rangle e_{i}$ and is of the form $g\delta^{a}e_{i}$ with $g\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$. By definition of $f$, the polynomial $g$ divides a power of $f$. let us denote $f^{n}$ this power, then
$\operatorname{in}_{(\bm{0},\bm{1})}\!\left(\frac{f^{n}}{g}L\right)=f^{n}\delta^{a}e_{i},$
@@
Finally, let $s\in\operatorname{Cl}_{\mathbf{x}}(S)$. There exists $q\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ such that $qa\in S$. Thus the class of $qa$ vanishes in $N$, and torsion-freeness implies that the class of $a$ also vanishes. Hence $a\in S[1/f]$. Since $a\in W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$, this proves the result. ∎
@@
(16) $f^{j+1}\partial_{x_{i}}L=\left(\partial_{x_{i}}f+(j+1)\frac{\partial f}{\partial x_{i}}\right)f^{j}L.$
@@
Let $S$ be a submodule of $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ and let $f\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ be a polynomial vanishing on $\operatorname{Sing}(S)$. Then,
@@
The ideal generated by $g_{1},g_{2},g_{3}$ corresponds the Weyl closure of $W_{x,y}g_{1}+W_{x,y}g_{2}$, which is also the annihilator of $f$ in $W_{x,y}$.
@@ 5. The algorithm
$\partial_{\ell}(Tf-1)=(Tf-1)(\partial_{\ell}+Tf_{\ell})$
for $\ell\in\{\mathbf{x},\mathbf{t}\}$ and the reverse inclusion follows from the analog formula
$(Tf-1)\partial_{\ell}=(\partial_{\ell}-Tf_{\ell})(Tf-1).$
@@
Besides fT1fT-1 divides $T^{i}f^{i}-1$, which proves by the remark before Lemma 2 that $(T^{i}f^{i}-1)$ is in $U_{f}$. This proves that $L$ is in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]S+U_{f}$. Because $L$ is in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})$ by definition, it belongs to the intersection.
@@
Lemma 1 suggests a strategy for computing the partial Weyl closure using saturation. In the commutative case this would be done by fixing a monomial order eliminating $T$ and computing a Gröbner basis with respect to this order. This is however not possible in this non-commutative setting as there is no order on $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ eliminating $T$ that is compatible with multiplication, in the sense that the product of leading terms is the leading term of the product (recall Eq. 20). This problem is tackled by seeing $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ as an infinite-rank $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})$-module and performing Gröbner basis computation on truncations of this infinite rank module. This idea of computing Gröbner bases of truncations of an infinite-rank module has already been used by Takayama in his integration algorithm (Takayama, 1990). As in his original algorithm, the termination criterion used here is based on holonomicity and does not guarantee that the full saturation module is obtained.

### v2

# Computing a holonomic submodule of the partial Weyl closure

@@ Abstract.
The Weyl closure is a basic operation in algebraic analysis: it converts a system of differential operators with rational coefficients into an equivalent system with polynomial coefficients. In addition to encoding finer information on the singularities of the system, it serves as a preparatory step for many algorithms in symbolic integration. A new algorithm is introduced to compute a holonomic submodule of the partial Weyl closure of a finite-rank module, where the closure is taken with respect to a subset of the variables. The method relies on a non-commutative analogue of Rabinowitsch’s trick. The algorithm is implemented in the Julia package MultivariateCreativeTelescoping.jl and shows substantial speedups over existing exact Weyl closure algorithms in Singular and Macaulay2.
@@ 1. Introduction
A recent integration algorithm (Brochet et al., 2025) assumes that the parameters $\mathbf{t}=(t_{1},\dots,t_{m})$ of the integrals are rational variables, while the integration variables $\mathbf{x}$ are polynomial variables, and takes as input generators of $\operatorname{ann}_{W_{t,\mathbf{x}}(\mathbf{t})}(f)$ where $W_{t,\mathbf{x}}(\mathbf{t})=\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\partial_{\mathbf{t}},\partial_{\mathbf{x}}\rangle$. This motivates extending the definition of Weyl closure to partial Weyl closure. Let $S\subseteq W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}$ be a left submodule. The partial Weyl closure of $S$ with respect to $\mathbf{x}$ is
@@
The problem of computing generators of $\operatorname{Cl}(S)$ given $S\subset W_{\mathbf{x}}^{r}$ was first introduced by Chyzak and Salvy (Chyzak and Salvy, 1998) for the case $r=1$, under the name of extension/contraction problem. This problem was later solved by Tsai for general $r$ in his PhD thesis (Tsai, 2000b, a). Tsai showed (Tsai, 2000a, Theorem 2.2.1) that the Weyl closure of $S\subset W_{\mathbf{x}}^{r}$ can be computed by inverting a single polynomial $p$ vanishing on the singular locus of $S$:
@@
The computation of annihilators of holonomic functions in the rational Weyl algebra $W_{\mathbf{x}}(\mathbf{x})$ is already well understood. Systems of differential equations are known for basic functions and algorithmic closure by sum, product and more (Comtet, 1964; Takayama, 1992) exist for deducing systems for more complicated ones. These algorithms are implemented in Maple (Chyzak, ), Mathematica (Koutschan, 2014), and Sage (Kauers et al., 2015). Although similar closure properties for sum and product exist in the Weyl algebra (Oaku and Takayama, 2001) (see also (Oaku, 2013, Sec.3)), their computational efficiency seems unclear. For example, the closure by product requires doubling the number of variables and eliminate half of them by Gröbner basis computation, which is known to be costly.
@@
The notion of holonomicity is extended to $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-modules and a new algorithm is proposed for computing a holonomic submodule of the partial Weyl closure. That is, given a finite rank module $S\subset W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}$, the algorithm computes a module $S^{\prime}$ such that $S^{\prime}\subseteq\operatorname{Cl}(S)$ and $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}/S^{\prime}$ is holonomic. The algorithm proceeds by iteratively constructing a sequence of modules
@@
The algorithm is based on a generalization of Rabinowitsch’s trick to D-modules, which consists in introducing a new variable $T$ thought of as $1/p$ for some $p\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$, together with the additional equation pT1=0pT-1=0. The commutation rules between $T$ and $\partial_{\mathbf{x}},\partial_{\mathbf{t}}$ are not well suited for computing Gröbner basis in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ for an order eliminating $T$. This problem is tackled by computing Gröbner bases in $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})[T]^{r}$ seen as an infinite-rank $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-module (Takayama, 1990). Then, saturations of $S$ with respect to $p$ are computed iteratively until the stopping condition is reached. Lastly, the algorithm has been implemented in the Julia package MultivariateCreativeTelescoping.jl (Brochet, 2025b). It compares very favorably to existing implementations.
Compared with Tsai’s approach, the present algorithm avoids several computationally expensive steps. Tsai’s algorithm first computes generators of the $W_{\mathbf{x}}$-module $W_{\mathbf{x}}[1/p]S$ using the localization algorithm of Oaku, Takayama, and Walther (Oaku et al., 2000), and then recovers the Weyl closure by computing the intersection $W_{\mathbf{x}}[1/p]S\cap W_{\mathbf{x}}^{r}$. The localization step requires the computation of a $b$-function associated with a weight vector of a suitable ideal, which is known to be difficult in general. Moreover, simplifications with respect to the polynomial $p$ are postponed until the final intersection step. In particular, when $S$ is an ideal annihilating a function $f$, the localization algorithm returns an ideal $S^{\prime}$ together with an integer $\ell$ such that $S^{\prime}$ annihilates $f/p^{\ell}$, which typically leads to a significant blow-up in the size of expressions compared to $S$ and $\operatorname{Cl}(S)$. This issue is avoided in the present algorithm by performing simplifications with respect to $p$ as early as possible through iterative saturation.
@@ 2.2. Singular Locus
The principal symbol (or initial form) of $P$ with respect to $(\bm{0},\bm{1})$ is the element in $\mathbb{K}(t)[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$ defined by
@@
$\operatorname{in}_{(\bm{0},\bm{1})}(S)\subset\mathbb{K}(t)[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$
@@
$\operatorname{char}(M)=V\!\left(\operatorname{in}_{(\bm{0},\bm{1})}(S)\right)\;\subset\;\mathbb{K}(t)^{2n+m}.$
@@
$\mathbb{K}(t)^{2n+m}\to\mathbb{K}(t)^{n},\qquad(\mathbf{x},{\bm{\xi}},{\bm{\zeta}})\mapsto\mathbf{x}.$
@@
$\operatorname{Sing}(I)=V\left(\left(\operatorname{in}_{(\bm{0},\bm{1})}(S):\langle\xi_{1},\dots,\xi_{n},\zeta_{1},\dots,\zeta_{m}\rangle^{\infty}\right)\cap\mathbb{K}(t)[\mathbf{x}]\right),$
where, for a module $A\subset\mathbb{K}(t)[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}$, and an ideal $B\subset\mathbb{K}(t)[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]$ the saturation of $A$ with respect to $B$ is given by
(10) $A:B^{\infty}=\left\{a\in\mathbb{K}(t)[\mathbf{x},{\bm{\xi}},{\bm{\zeta}}]^{r}\;\middle|\;\exists\,k\in\mathbb{N}\text{ such that }B^{k}a\subset A\right\}.$
@@ 3. Holonomic $W_{\mathbf{t},\mathbf{x}}(\mathbf{t})$-modules
When $m=0$, meaning that no variables $t_{i}$ are present, this definition agrees with the standard definition of holonomic $W_{\mathbf{x}}$-modules by Theorem 1, and when $n=0$, meaning that no variables $x_{i}$ are present, it agrees with the definition of D-finite modules, that is, modules that are finite-dimensional over $\mathbb{K}(\mathbf{t})$.
The equivalence between the second and third point of Theorem 1, which gave an algorithmic criterion for testing holonomicity, naturally extends to this setting.
@@ 4. Partial Weyl closure by saturation
The goal of this section is to prove Theorem 1, which states that the partial Weyl closure of a module $S\subset W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ can be computed by inverting a single polynomial vanishing on the singular locus of $S$. A similar result was proved by Tsai (Tsai, 2000a, Theorem 2.2.1) in the case where no parameters $\mathbf{t}$ is present. The argument given here is a straightforward extension of his approach to this mixed setting.
@@
Let $S$ be a submodule of $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ with finite rank and let $f\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ be a polynomial vanishing on $\operatorname{Sing}(S)$. Then,
@@
Proof of Lemma 2.
Let $T$ be the torsion module
@@
annihilates every element of $T$. Applying any $\delta\in\{\bm{\partial}_{\mathbf{x}},\bm{\partial}_{\mathbf{t}}\}$ for each $m\in T$ gives
@@
Indeed, since $S$ is a finite-rank module, for any $i\in\llbracket 1,r\rrbracket$ and any $\varsigma\in\{\bm{\partial}_{\mathbf{x}},\bm{\partial}_{\mathbf{t}}\}$, there exists a nonzero element $L^{\prime}\in W_{\mathbf{t},\mathbf{x}}(\mathbf{t},\mathbf{x})S\cap\mathbb{K}(\mathbf{t},\mathbf{x})\langle\varsigma\rangle e_{i}$. Multiplying it by a suitable polynomial, we obtain a non-zero element $L\in S\cap\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\varsigma\rangle e_{i}$. Its initial form $\operatorname{in}_{(\bm{0},\bm{1})}(L)$ then belongs to $\mathbb{K}(\mathbf{t})[\mathbf{x}]\langle\varsigma\rangle e_{i}$ and is of the form $g\varsigma^{a}e_{i}$ with $g\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$. By definition of $f$, the polynomial $g$ divides a power of $f$. let $f^{n}$ denote this power, then
$\operatorname{in}_{(\bm{0},\bm{1})}\!\left(\frac{f^{n}}{g}L\right)=f^{n}\varsigma^{a}e_{i},$
@@
Finally, let $a\in\operatorname{Cl}_{\mathbf{x}}(S)$. There exists $q\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ such that $qa\in S$. Thus the class of $qa$ vanishes in $N$, and torsion-freeness implies that the class of $a$ also vanishes. Hence $a\in S[1/f]$. Since $a\in W_{\mathbf{t},\mathbf{x}}(\mathbf{t})^{r}$, this proves the result. ∎
@@
(16) $f^{j+1}\partial_{x_{i}}L=\left(\partial_{x_{i}}f-(j+1)\frac{\partial f}{\partial x_{i}}\right)f^{j}L.$
@@
Let $S$ be a submodule of $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})^{r}$ with finite rank and let $f\in\mathbb{K}(\mathbf{t})[\mathbf{x}]$ be a polynomial vanishing on $\operatorname{Sing}(S)$. Then,
@@
The ideal generated by $g_{1},g_{2},g_{3}$ corresponds to the Weyl closure of $W_{x,y}g_{1}+W_{x,y}g_{2}$, which is also the annihilator of $f$ in $W_{x,y}$.
@@ 5. The algorithm
$\partial_{\ell}(Tf-1)=(Tf-1)(\partial_{\ell}-Tf_{\ell})$
for $\ell\in\{\mathbf{x},\mathbf{t}\}$ and the reverse inclusion follows from the analogous formula
$(Tf-1)\partial_{\ell}=(\partial_{\ell}+Tf_{\ell})(Tf-1).$
@@
Besides fT1fT-1 divides $T^{i}f^{i}-1$, which implies by Lemma 2 that $(T^{i}f^{i}-1)$ is in $U_{f}$. This proves that $L$ is in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]S+U_{f}$. Because $L$ is in $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})$ by definition, it belongs to the intersection.
@@
Lemma 1 suggests a strategy for computing the partial Weyl closure using saturation. In the commutative case this would be done by fixing a monomial order eliminating $T$ and computing a Gröbner basis with respect to this order. However, this is not possible in this non-commutative setting as there is no order on $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ eliminating $T$ that is compatible with multiplication, in the sense that the product of leading terms is the leading term of the product (recall Eq. 20). This problem is tackled by seeing $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})[T]^{r}$ as an infinite-rank $W_{\mathbf{x},\mathbf{t}}(\mathbf{t})$-module and performing Gröbner basis computation on truncations of this infinite-rank module. This idea of computing Gröbner bases of truncations of an infinite-rank module has already been used by Takayama in his integration algorithm (Takayama, 1990). As in his original algorithm, the termination criterion used here is based on holonomicity and does not guarantee that the full saturation module is obtained.